OpenStudy (anonymous):

differentiate: y(t)=(1)/(t+4e^t)

OpenStudy (anonymous):

so i know well be using the product rule..but i just dont know what to do with it after i took the derivative

OpenStudy (nice):

it's chain rule : y= (t + 4e^t)^-1 dy/dt = [-(t + 4e^t)^-2] x derivative of (t + 4e^t) dy/dt = [-(t + 4e^t)^-2] x (1 + 4e^t)

OpenStudy (anonymous):

its a fraction 1 over t + 4e^t

OpenStudy (nice):

\[[-(t + 4e^t)^-2] \times d(t + 4e^t)/dt\] \[dy/dt = [-(t + 4e^t)^-2] \times (1 + 4e^t)\]

OpenStudy (nice):

aha .. so you differentiate the outer as it is u^-1 ,, then multiply it by the derivative of the inner (t + 4e^t) this is the chain rule yes I make 1 over t + 4e^t = (t + 4e^t)^-1

OpenStudy (nice):

did you understand ??

OpenStudy (anonymous):

\[d/dx{\frac{1}{f(x)}=\frac{-f'(x)}{f^2(x)}}\] so you answer here is \[\frac{-(4e^x+1)}{(x+4e^x)^2}\]